OpenStudy (anonymous):

Solve by simulating the problem. You have a 5-question multiple-choice test. Each question has four choices. You don’t know any of the answers. What is the experimental probability that you will guess exactly three out of five questions correctly? Type your answer below using complete sentences.

OpenStudy (anonymous):

lol i googled this 45/512

OpenStudy (anonymous):

heres the link

OpenStudy (anonymous):

i have to kinda explain it

OpenStudy (anonymous):

i dont know how to do this bro, thats why i sent you the link.

OpenStudy (anonymous):

20 answer choices and only 5 are right would amount to threoretical prob

OpenStudy (anonymous):

Hahah okay well thanks!

OpenStudy (anonymous):

thank you @Gio123

OpenStudy (anonymous):

your welcome. but that was only the "theoretical probability" there is no way to get any other unless you actually take a test and simulate it like that. But yeah, theoretically that is the answer. :) hoped i helped.

OpenStudy (ybarrap):

Using combination notation: $$ \large \bf \binom{5}{3}\dfrac{1}{4}^3\dfrac{3}{4}^2 =\dfrac{5\times4}{2}\left(\dfrac{1}{4}\right)^3\left(\dfrac{3}{4}\right)^2=\dfrac{45}{512} $$